# MagicColor: Multi-Instance Sketch Colorization

### v1

@@ Abstract
We present MagicColor, a diffusion-based framework for multi-instance sketch colorization. The production of multi-instance 2D line art colorization adheres to an industry-standard workflow, which consists of three crucial stages: the design of line art characters, the coloring of individual objects, and the refinement process. The artists are required to repeat the process of coloring each instance one by one, which is inaccurate and inefficient. Meanwhile, current generative methods fail to solve this task due to the challenge of multi-instance pair data collection. To tackle these challenges, we incorporate three technical designs to ensure precise character detail transcription and achieve multi-instance sketch colorization in a single forward. Specifically, we first propose the self-play training strategy to solve the lack of training data. Then we introduce an instance guider to feed the color of the instance. To achieve accurate color matching, we present fine-grained color matching with edge loss to enhance visual quality. Equipped with the proposed modules, MagicColor enables automatically transforming sketches into vividly-colored images with accurate consistency and multi-instance control. Experiments on our collected datasets show that our model outperforms existing methods regarding chromatic precision. Specifically, our model critically automates the colorization process with zero manual adjustments, so novice users can produce stylistically consistent artwork by providing reference instances and the original line art. Our code and additional details are available at https://yinhan-zhang.github.io/color.
Figure 1: Visual results of MagicColor. Given a set of references and corresponding line art images, our approach enables coloring sketches while maintaining consistency across multiple instances. Compared to traditional methods, our approach significantly improves coloring efficiency.
@@
Figure 2: Illustration of the workflow of multi-instance sketch colorization production. Previous methods can only achieve multi-instance sketch colorization step by step, which is time-consuming and inaccurate. In contrast, our method can color sketch while maintaining consistency, making multi-instance sketch colorization easier.
@@ 1 Introduction
The fast development of the digital cartoon industry has demonstrated widespread potential applications for generative artificial intelligence. Colorizing cartoon sketches, a crucial aspect within the broader realm of the cartoon industry, serves as an indispensable task. It not only heightens the visual allure but also enriches the narrative experience by effectively vividly communicating emotions and actions. Using the automatic pipeline can streamline production workflows, accelerate content creation, reduce labor demands, and meet the growing demand for cartoons.
In the traditional process of cartoon sketch colorization, artists first analyze the line art to grasp the character and story context. They then choose suitable color palettes and manually apply colors layer by layer with attention to detail. However, traditional methods have two drawbacks. The first one is that they are time-consuming as manual coloring requires great effort to match the original design. For instance, in an animation series, multiple identical objects in the same image need to be colored one by one, which is very inefficient. The other problem is inaccuracy since manual work is error-prone and different artists may have varied color interpretations, causing inconsistent results in large-scale projects. Our method slots into this workflow, supporting auto-colorization while preserving design fidelity and instance-level color consistency.
@@
To overcome these challenges, we propose MagicColor, a novel framework that streamlines the colorization process. Our approach builds upon pre-trained diffusion model priors, capitalizing on their learned capacity to enforce visual consistency across generated outputs. The architectural cornerstones of our approach are articulated as follows. First, to tackle the misalignment between the reference instances and target line art, we incorporate an explicit reference mechanism that seamlessly infuses reference-derived color semantics and artistic styles into line art. In addition, we employ implicit latent control to ensure precise reference to each instance, significantly enhancing color accuracy and consistency. Second, to further improve the visual quality and color consistency of the output, we introduce edge loss and color matching. These methods compel the model to genuinely extract color information from the reference character design, thereby improving the accuracy of semantic correspondence and reducing the reliance on any color information that may inadvertently leak from the reference image. Third, our model adopts a two-stage, self-play training strategy, which addresses the challenges of limited multi-instance training data, and subsequently incorporates additional reference images to refine the colorization capability. By facilitating colorization across multiple instances, our model achieves impressive color consistency with minimal human intervention. Our approach achieves state-of-the-art results across both quantitative metrics and qualitative evaluations, outperforming prior art in animated content creation. We aim for a critical step toward fully automated, high-efficiency animation pipelines with guaranteed stylistic coherence. This method can also be extended beyond anime to the broader digital art and media fields. Our contributions can be summarized as follows:
  • We propose MagicColor, the first multi-instance coloring method to support multiple instances integration for sketch colorization in a single forward.
  • Technically, to solve the lack of multi-instance data, we design a two-stage, self-play training strategy. We also propose instance guider and pixel-level color matching with edge loss to enhance the color correspondence.
@@ 2.2 Visual Correspondence
In computer vision, visual correspondence aims to identify and match relevant features or points among different images. It is widely applied in tasks such as stereo vision and motion tracking. In the past, traditional methods [2, 46] relied on hand-designed features to establish corresponding relationships. Nowadays, deep learning methods [12, 31, 28, 9] obtain matching capabilities through supervised learning with labeled datasets. However, supervised learning faces significant scalability issues. Precise pixel-level annotations are not only time-consuming and labor-intensive but also costly. To address this, scholars have started to explore weakly-supervised or self-supervised visual correspondence models. For example, LightGlue [38] can match sparse local features across image pairs, through an adaptive mechanism. CoTracker [24] adopts a semi-supervised training method, by generating pseudo-labels using off-the-shelf models. DIFT [74] extracts features by diffusion models and can achieve pixel-level semantic point matching. Building upon the diffusion models’ prior correspondence knowledge, our framework enables reference-based colorization via semantic-aligned color mapping between line art and reference instances, without any structural modifications.
@@ 2.3 Reference-Based Image Colorization
A great deal of research has been committed to colorizing photographs using reference-based priors [18, 4, 91, 15, 75]. Initially, efforts focused on the transfer of chromatic information to the corresponding regions through luminance and texture alignment, with various low-level feature-based correspondence techniques developed for more precise local color transfer [32, 41, 7, 80, 53]. However, these methods are vulnerable to complex appearance variations of the same object as low-level features cannot capture semantic nuances [17, 99]. Line art colorization is notably different from natural image colorization [66, 84, 96, 89].
@@ 3.1 Latent Diffusion Model
As the core architecture of Stable Diffusion [61], Latent Diffusion Models (LDM) revolutionize text-to-image generation by executing diffusion-denoising processes in compressed latent space rather than raw pixel domain, enabling stable and efficient training. The pipeline begins with a Variational Autoencoder (VAE) projecting RGB images into low-dimensional latent codes, where semantic-guided diffusion sampling occurs under textual conditioning. Then, an UNet-based network incorporates self-attention and cross-attention mechanisms through UNet blocks to learn the reverse denoising process in latent space. Cross-attention establishes bidirectional interactions between text embeddings and visual features, ensuring prompt semantics are continuously infused. The whole training objective of the UNet can be written as:
@@
where $z$ notes the latent embedding of the training sample. $\epsilon_{\theta}$ and $\epsilon$ represent predicted noise by diffusion model and ground truth noise at corresponding timestep $t$, respectively. $c$ is the condition embedding involved in the generation and the co-efficient $\alpha_{t}$ remains consistent with that employed in vanilla diffusion models.
@@ 3.2 Reference Condition Injection
In the current scenario, reference instances frequently suffer from noise redundancy and semantic conflicts, which significantly impede the model’s ability to learn colors effectively. To address this issue, when presented with $N$ user-provided reference instances along with the line art input, our model employs a dual-branch condition injection strategy. This strategy aims to achieve semantic similarity and structural alignment with the input. First, we align multiple instances in a layer and input them into the reference net. Then, we encode with CLIP and apply reference attention mechanisms [18, 4, 45] to inject color and semantic info into the UNet. Also, the line art is injected into the UNet to enhance line-structure features. This lets the model learn global reference info evenly and avoid interference. Second, for instance-aware image generation, we utilize instance images $I_{i}$ embedding as latent control signals. Unlike other methods [63, 45], our paradigm supports the use of multiple instances, which enhances the model’s generalizability. Zero-shot customization is tough, so we use pre-trained vision models to extract target object identity. Previous studies used CLIP for target embedding [34, 60]. We use DINOv2 as the feature encoder [55] to get discriminative spatial identity. DINOv2, trained with patch-level objectives under random masking, has highly expressive features. Its output includes a $26\times 26\times 1024$ spatial embedding $s_{i}$ for patch-level features and a $1024$ dimensional global embedding $g_{i}$.
Figure 3: The overview of MagicColor pipeline. We combine a dual-UNet framework with an instance control module (ICM). During training, we use multiple instances to set the overall color accurately. The color matching enables the model to better align the colors of the target image with those of the reference instances precisely. The edge loss helps the model pay more attention to the high-frequency areas and edges, resulting in a more accurate and vivid colorization for each instance.
@@ 4 Method
where $I_{out}$ is the output image to be generated. The condition $C$ is composed of a line art image $S$, a set of reference instances $R=\{[I_{1},\cdots,I_{i}]\mid i=1,\cdots,N\}$, and a set of instance masks $M=\{[M_{1},\cdots,M_{i}]\mid i=1,\cdots,N\}$. Each $I_{i}$ in $R$ corresponds to an instance, and each $M_{i}$ in $M$ is a binary mask indicating the reference spatial location of an instance.
Architecture Design. Owing to the high-fidelity detail demands in anime sketch colorization, the core challenge resides in designing encoders capable of spatially precise reference image parsing to extract sub-pixel-level visual cues. Inspired by recent studies [19, 35, 45] which demonstrated the effectiveness of leveraging an additional UNet architecture for this purpose, we introduce a reference net following a similar design. Additionally, the sketch and instance guider are designed based on ControlNet [97], which provides an efficient condition for injecting line art and instance-level information into the generative process. Furthermore, we employ DINOv2 [55] to encode images and train a Feed-Forward Network layer to extract features from the image embeddings. These extracted features are then integrated into the latent control signals.
@@ 4.1 Self-Play Training Strategy
We then introduce a two-stage, self-play training approach to solve the problem of lack of multi-instance training data and gradually boost the model’s performance to build an advanced model that can give high-quality coloring results.
Single-Reference Colorization Training. In the first stage, training starts by activating the reference net, UNet, and the sketch guider. In both two stages, we use color matching with edge loss for training. For each anime sequence, we perform random frame sampling without replacement: one frame is designated as the style reference exemplar, while the other serves as the raw input sketch for colorization. We use the whole reference image as conditional input and extract the line art from the original image. Given a reference and line art image pair, the model maps color semantics from the reference image to the line art by minimizing a carefully designed edge loss function, aiming to make the model learn basic color-semantic relationships, forming the basis for later multi-instance refinement.
@@ 4.2 Instance Control Module
Previous methods often resorted to directly imposing constraints within the reference attention mechanism for control purposes [18, 4, 19]. Nevertheless, when faced with the task of handling multiple instances as conditional input, this conventional approach proved to be extremely arduous. To uphold semantic correspondence and safeguard details, we align image embeddings in latent space, thereby empowering more precise and effective handling of the intricate input scenarios presented by multiple instances. Given target regions (instance masks), we align the bounding rectangle of the mask with the input DINO feature maps using box coordinates. Subsequently, we use the dense grids of the mask within this rectangle to interpolate features from the input DINO features. This process is similar to ROI alignment, but the resolution of ROI is flexible and follows the size of the mask. Finally, to further incorporate global information, we randomly drop 10% of the spatial embedding $s_{i}$ and replace it with the DINO global embedding $g_{i}$. As our target mask may have a different silhouette from the shape of the input image, injecting global embedding enhances the model’s generalizability and adaptability across diverse image conditions. The final ROI feature is then warped into the original region of the latent control signal:
@@ 4.3 Structure-Content Supervision Enhancement
Figure 5: Inference Pipeline. When users input a line art alongside one or multiple reference instances, our model automatically fuses instances across different layers, enhances line trajectories in the line art, and then performs colorization.
Color Matching. We employ pre-trained diffusion models to identify corresponding points within real-world images. Our model, which has a UNet at its core, processes noisy images by cleaning them and extracting features crucial for establishing correspondences. Subsequently, we utilize a diffusion feature map to match pixels between two images. This is accomplished by determining the nearest neighbors and computing the similarity using cosine distance. To obtain pixel correspondences, we begin by extracting dense features from both images and matching them. Let $F$ be the feature map of an image $I$. For a pixel at position $p$, we obtain the feature vector $F(p)$ through bilinear interpolation. In terms of color matching, $C_{r}$ denotes the color feature of the reference image, and $C_{s}$ denotes the color feature of the source image. We obtain $C_{s}$ and $C_{r}$ by sampling features from the normalized source and target features. To quantify their similarity, we calculate the Euclidean distance ($D$) between $C_{r}$ and the flattened $C_{s}$. Given two feature vectors $v_{1}=C_{r}$ and $v_{2}=C_{s}$, the Euclidean distance is computed as follows:
@@
where $f$ represents the feature dimension. Subsequently, we find the nearest-neighbor indices. Finally, given a pixel $p_{1}$ in $v_{1}$, we find the corresponding pixel $p_{2}$ in $v_{2}$ as follows:

### v2

# Follow-Your-Color: Multi-Instance Sketch Colorization

@@ Abstract
We present Follow-Your-Color, a diffusion-based framework for multi-instance sketch colorization. The production of multi-instance 2D line art colorization adheres to an industry-standard workflow, which consists of three crucial stages: the design of line art characters, the coloring of individual objects, and the refinement process. The artists are required to repeat the process of coloring each instance one by one, which is inaccurate and inefficient. Meanwhile, current generative methods fail to solve this task due to the challenge of multi-instance pair data collection. To tackle these challenges, we incorporate three technical designs to ensure precise character detail transcription and achieve multi-instance sketch colorization in a single forward pass. Specifically, we first propose the self-play training strategy to address the lack of training data. Then we introduce an instance guider to feed the color of the instance. To achieve accurate color matching, we present fine-grained color matching with edge loss to enhance visual quality. Equipped with the proposed modules, Follow-Your-Color enables automatically transforming sketches into vividly-colored images with accurate consistency and multi-instance control. Experiments on our collected datasets show that our model outperforms existing methods regarding chromatic precision. Specifically, our model critically automates the colorization process with zero manual adjustments, so novice users can produce stylistically consistent artwork by providing reference instances and the original line art. Our code and additional details are available at https://yinhan-zhang.github.io/color.
Figure 1: Results of Follow-Your-Color. Given a set of colored references, Follow-Your-Color can colorize a line art image while maintaining color consistency across multiple instances. Compared to traditional methods, our approach significantly improves coloring efficiency.
@@
Figure 2: Illustration of the workflow of multi-instance sketch colorization production. Previous methods can only achieve multi-instance sketch colorization step by step, which is time-consuming and inaccurate. In contrast, our method can color a sketch while maintaining consistency, making multi-instance sketch colorization easier.
@@ 1 Introduction
The fast development of the digital cartoon industry has demonstrated widespread potential applications for generative artificial intelligence. Colorizing cartoon sketches, a crucial aspect within the broader realm of the cartoon industry, serves as an indispensable task. It not only heightens the visual allure but also enriches the narrative experience by effectively and vividly communicating emotions and actions. Using the automatic pipeline can streamline production workflows, accelerate content creation, reduce labor demands, and meet the growing demand for cartoons.
In the traditional process of cartoon sketch colorization, artists first analyze the line art to grasp the character and story context. They then choose suitable color palettes and manually apply colors layer by layer with attention to detail. However, traditional methods have two drawbacks. The first one is that they are time-consuming, as manual coloring requires great effort to match the original design. For instance, in an animation series, multiple identical objects in the same image need to be colored one by one, which is very inefficient. The other problem is inaccuracy since manual work is error-prone and different artists may have varied color interpretations, causing inconsistent results in large-scale projects. Our method slots into this workflow, supporting auto-colorization while preserving design fidelity and instance-level color consistency.
@@
To overcome these challenges, we propose Follow-Your-Color, a novel framework that streamlines the colorization process. Our approach builds upon pre-trained diffusion model priors, capitalizing on their learned capacity to enforce visual consistency across generated outputs. The architectural cornerstones of our approach are articulated as follows. First, to tackle the misalignment between the reference instances and target line art, we incorporate an explicit reference mechanism that seamlessly infuses reference-derived color semantics and artistic styles into line art. In addition, we employ implicit latent control to ensure precise reference to each instance, significantly enhancing color accuracy and consistency. Second, to further improve the visual quality and color consistency of the output, we introduce edge loss and color matching. These methods compel the model to genuinely extract color information from the reference character design, thereby improving the accuracy of semantic correspondence and reducing the reliance on any color information that may inadvertently leak from the reference image. Third, our model adopts a two-stage, self-play training strategy, which addresses the challenges of limited multi-instance training data and subsequently incorporates additional reference images to refine the colorization capability. By facilitating colorization across multiple instances, our model achieves impressive color consistency with minimal human intervention. Our approach achieves state-of-the-art results across both quantitative metrics and qualitative evaluations, outperforming prior art in animated content creation. We aim for a critical step toward fully automated, high-efficiency animation pipelines with guaranteed stylistic coherence. This method can also be extended beyond anime to the broader digital art and media fields. Our contributions can be summarized as follows:
  • We propose Follow-Your-Color, the first multi-instance coloring method to support multiple instances integration for sketch colorization in a single forward pass.
  • Technically, to solve the lack of multi-instance data, we design a two-stage, self-play training strategy. We also propose an instance guider and pixel-level color matching with edge loss to enhance the color correspondence.
@@ 2.2 Visual Correspondence
In computer vision, visual correspondence aims to identify and match relevant features or points among different images. It is widely applied in tasks such as stereo vision and motion tracking. In the past, traditional methods [2, 45] relied on hand-designed features to establish corresponding relationships. Nowadays, deep learning methods [12, 31, 28, 9] obtain matching capabilities through supervised learning with labeled datasets. However, supervised learning faces significant scalability issues. Precise pixel-level annotations are not only time-consuming and labor-intensive but also costly. To address this, scholars have started to explore weakly-supervised or self-supervised visual correspondence models. For example, LightGlue [38] can match sparse local features across image pairs through an adaptive mechanism. CoTracker [24] adopts a semi-supervised training method by generating pseudo-labels using off-the-shelf models. DIFT [73] extracts features by diffusion models and can achieve pixel-level semantic point matching. Building upon the diffusion models’ prior correspondence knowledge, our framework enables reference-based colorization via semantic-aligned color mapping between line art and reference instances, without any structural modifications.
@@ 2.3 Reference-Based Image Colorization
A great deal of research has been committed to colorizing photographs using reference-based priors [18, 4, 90, 15, 74]. Initially, efforts focused on the transfer of chromatic information to the corresponding regions through luminance and texture alignment, with various low-level feature-based correspondence techniques developed for more precise local color transfer [32, 40, 7, 79, 52]. However, these methods are vulnerable to complex appearance variations of the same object, as low-level features cannot capture semantic nuances [17, 98]. Line art colorization is notably different from natural image colorization [65, 83, 95, 88].
@@ 3.1 Latent Diffusion Model
As the core architecture of Stable Diffusion [60], Latent Diffusion Models (LDM) revolutionize text-to-image generation by executing diffusion-denoising processes in a compressed latent space rather than the raw pixel domain, enabling stable and efficient training. The pipeline begins with a Variational Autoencoder (VAE) projecting RGB images into low-dimensional latent codes, where semantic-guided diffusion sampling occurs under textual conditioning. Then, a UNet-based network incorporates self-attention and cross-attention mechanisms through UNet blocks to learn the reverse denoising process in the latent space. Cross-attention establishes bidirectional interactions between text embeddings and visual features, ensuring prompt semantics are continuously infused. The whole training objective of the UNet can be written as:
@@
where $z$ notes the latent embedding of the training sample. $\epsilon_{\theta}$ and $\epsilon$ represent predicted noise by the diffusion model and ground truth noise at corresponding timestep $t$, respectively. $c$ is the condition embedding involved in the generation, and the coefficient $\alpha_{t}$ remains consistent with that employed in vanilla diffusion models.
@@ 3.2 Reference Condition Injection
In the current scenario, reference instances frequently suffer from background noise, redundancy, and semantic conflicts (e.g., similar instances coexist in the same reference image, or the background contains complex textures), which significantly impede the model’s ability to learn colors effectively. To address this issue, when presented with $N$ user-provided reference instances along with the line art input, our model employs a dual-branch condition injection strategy. This strategy aims to achieve semantic similarity and structural alignment with the input. First, we align multiple instances in a layer and input them into the reference net. Then, we encode with CLIP and apply reference attention mechanisms [18, 4, 44] to inject color and semantic info into the UNet. Also, the line art is injected into the UNet to enhance line-structure features. This lets the model learn global reference info evenly and avoid interference. Second, for instance-aware image generation, we utilize instance images $I_{i}$ embedding as latent control signals. Unlike other methods [62, 44], our paradigm supports the use of multiple instances, which enhances the model’s generalizability. Zero-shot customization is tough, so we use pre-trained vision models to extract the target object identity. Previous studies used CLIP for target embedding [34, 59]. We use DINOv2 as the feature encoder [54] to get discriminative spatial identity. DINOv2, trained with patch-level objectives under random masking, has highly expressive features. Its output includes a $26\times 26\times 1024$ spatial embedding $s_{i}$ for patch-level features and a $1024$ dimensional global embedding $g_{i}$.
Figure 3: Overview of the Follow-Your-Color pipeline. We combine a dual-UNet framework with an instance control module (ICM). During training, we use multiple instances to set the overall color accurately. The color matching enables the model to better align the colors of the target image with those of the reference instances precisely. The edge loss helps the model pay more attention to the high-frequency areas and edges, resulting in a more accurate and vivid colorization for each instance.
@@ 4 Method
where $I_{out}$ is the output image to be generated. The condition is composed of a line art image $S$, a set of reference instances $R=\{[I_{1},\cdots,I_{i}]\mid i=1,\cdots,N\}$, and a set of instance masks $M=\{[M_{1},\cdots,M_{i}]\mid i=1,\cdots,N\}$. Each $I_{i}$ in $R$ corresponds to an instance, and each $M_{i}$ in $M$ is a binary mask indicating the reference spatial location of an instance.
Architecture Design. Owing to the high-fidelity detail demands in anime sketch colorization, the core challenge resides in designing encoders capable of spatially precise reference image parsing to extract pixel-level visual cues. Inspired by recent studies [19, 35, 44] which demonstrated the effectiveness of leveraging an additional UNet architecture for this purpose, we introduce a reference net following a similar design. Additionally, the sketch and instance guider are designed based on ControlNet [96], which provides an efficient condition for injecting line art and instance-level information into the generative process. Furthermore, we employ DINOv2 [54] to encode images and train a Feed-Forward Network layer to extract features from the image embeddings. These extracted features are then integrated into the latent control signals.
@@ 4.1 Self-Play Training Strategy
We then introduce a two-stage, self-play training approach to solve the problem of a lack of multi-instance training data and gradually boost the model’s performance to build an advanced model that can give high-quality coloring results.
Single-Reference Colorization Training. In the first stage, training starts by activating the reference net, UNet, and the sketch guider. In both stages, we use color matching with edge loss for training. For each anime sequence, we perform random frame sampling without replacement: one frame is designated as the style reference exemplar, while the other serves as the raw input sketch for colorization. We use the whole reference image as a conditional input and extract the line art from the original image. Given a reference and line art image pair, the model maps color semantics from the reference image to the line art by minimizing a carefully designed edge loss function, aiming to make the model learn basic color-semantic relationships, forming the basis for later multi-instance refinement.
@@ 4.2 Instance Control Module
Previous methods often resorted to directly imposing constraints within the reference attention mechanism for control purposes [18, 4, 19]. Nevertheless, when faced with the task of handling multiple instances as conditional input, this conventional approach proved to be extremely arduous. To uphold semantic correspondence and safeguard details, we align image embeddings in the latent space, thereby empowering more precise and effective handling of the intricate input scenarios presented by multiple instances. Given target regions (instance masks), we align the bounding rectangle of the mask with the input DINO feature maps using box coordinates. Subsequently, we use the dense grids of the mask within this rectangle to interpolate features from the input DINO features. This process is similar to ROI alignment, but the resolution of the ROI is flexible and follows the size of the mask. Finally, to further incorporate global information, we randomly drop 10% of the spatial embedding $s_{i}$ and replace it with the DINO global embedding $g_{i}$. As our target mask may have a different silhouette from the shape of the input image, injecting global embedding enhances the model’s generalizability and adaptability across diverse image conditions. The final ROI feature is then warped into the original region of the latent control signal:
@@
Table 3: Quantitative Comparisons Across Datasets
Domain Dataset FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Animation Hand-drawn Animation Hand-drawn Animation Hand-drawn Animation Hand-drawn
RSIC(256x) 128.971 300.636 10.933 8.872 0.548 0.416 0.508 0.566
SGA(256x) 124.732 306.881 10.021 7.253 0.521 0.414 0.454 0.533
AnimeDiffusion(256x) 154.784 305.021 11.262 10.207 0.416 0.426 0.557 0.549
ColorizeDiffusion(512x) 82.201 298.673 17.882 9.817 0.512 0.349 0.425 0.513
MangaNinja(512x) 43.165 295.223 14.289 10.249 0.543 0.361 0.362 0.508
Ours(512x) 28.953 251.301 23.751 10.551 0.783 0.434 0.201 0.431
@@ 4.3 Structure-Content Supervision Enhancement
Color Matching. We employ pre-trained diffusion models to identify corresponding points within real-world images. Our model, which has a UNet at its core, processes noisy images by cleaning them and extracting features crucial for establishing correspondences. Two steps utilize a diffusion feature map to match pixels between two images. (1) Semantic Matching. This is accomplished by determining the nearest neighbors and computing the similarity using cosine distance. To obtain pixel correspondences, we begin by extracting dense features from both images and matching them. Let $F_{i}$ be the feature map of an image $i$. For a pixel at position $p$, we obtain the feature vector $F(p)$ through bilinear interpolation. In terms of color matching, $C_{reference}$ denotes the color feature of the reference image, and $C_{source}$ denotes the color feature of the source image. We obtain $C_{source}$ and $C_{reference}$ by sampling features from the normalized source and target features. To quantify their similarity, we calculate the Euclidean distance ($D$) between $C_{reference}$ and the flattened $C_{source}$. Given two feature vectors $v_{1}=C_{reference}$ and $v_{2}=C_{source}$, the Euclidean distance is computed as follows:
@@
where $f$ represents the feature dimension. Subsequently, we find the nearest-neighbor indices. (2) Feature injection. Given a pixel $p_{1}$ in $v_{1}$, we find the corresponding pixel $p_{2}$ in $v_{2}$ as follows:
@@
Figure 5: Qualitative comparisons with existing methods. Given a line drawing and multiple reference instances, our method demonstrates far more precise colorization and higher-quality results compared to other methods, effectively maintaining line-drawing structure and reference instances’ color consistency.